OpenStudy (anonymous):

for f(x)=4x+1 and g(x)=x^2-5 find (g/f)(x)

OpenStudy (nevermind_justschool):

Welcome to Open Study!

OpenStudy (anonymous):

Thanks.

OpenStudy (deadly_roses):

try using wolframalpha.com it works really well and will show you how to do it, im not very good with these...or i would help...btw welcome to openstudy ^.^

OpenStudy (jhannybean):

\[\left(\frac{f}{g}\right)(x) = \frac{f(x)}{g(x)} = \frac{4x+1}{x^2-5}\]

OpenStudy (anonymous):

Thanks, both of you.

OpenStudy (jhannybean):

np.